solve the system by the substitution method.
15x-y=51
y=x^2+5

Answer by stanbon(75887)

solve the system by the substitution method.
15x-y=51
y=x^2+5
-------------
Substitute for "Y" and solve for "x":
15x-[x^2+5] = 51
15x-x^2-5 = 51
x^2-15x+56 = 0
Factor:
(x-7)(x-8) = 0
x = 7 or x = 8
----
Solve for "y":
y = x^2+5
If x = 7, y = 54
If x = 8, y = 69
=====================
Cheers,
Stan H.